A lead pellet is propelled vertically upward by an air rifle with an initial speed of 50 ft/s. Find its maximum height ?


1
Expert's answer
2022-02-13T12:16:18-0500

Given:

"v_0=50\:\rm ft/s"

"g=32.17\:\rm ft/s^2"


The maximum height


"h_{\max}=\frac{v_0^2}{2g}=\frac{50^2}{2*32.17}=39\:\rm ft"